Assume that Javier is indifferent between investing in a city of El Paso bond that pays 5 percent interest and a corporate bond that pays 6.25 percent interest. What is Javier's marginal tax rate?
Answer #1

Marginal tax rate is the rate at which an additional dollar of taxable income would be taxed. It is part of a progressive tax system, which applies different tax rates to different levels of income. As income rises, it is taxed at a higher rate (according to the bracket it falls in).

It’s important to note that income is not taxed at a single rate, but at multiple rates according to each tax bracket it falls in. Each tax rate only applies to income that falls within the corresponding bracket, and each additional dollar beyond that bracket would be taxed at the next highest marginal rate.

Interest Rate of El Paso Bond is 5%

Interest Rate of Corporate Bond is 6.25%

Marginal Tax Rate = Changes in the Tax Paid / Changes in Taxable Income

Suppose 1,00,000 is the Investment Made in EL Paso Bond, Income is 1,00,000 * 5/100 = 5,000

Suppose 1,00,000 is the Investment Made in Corporate Bond, Income is 1,00,000 * 6.25/100 = 6,250

Difference in the Taxable Incomes = 6,250-5,000 = 1,250